OpenStudy (kaleineedshelp):

Please help me with this, I recently started Algebra 1 and I'm already stuck and I really don't want to fail so here's the problem: 1/3^2 • 1/3^3 It looks like such a simple problem. I have answers in the back of my book but I want to know how to do it because my teachers gives daily quizzes on each topic we cover so please help me out. ( • = multiplication ) Thanks in advance.

OpenStudy (ememnim):

is that one over three?

OpenStudy (sshayer):

\[\frac{ 1 }{ 3*3 }*\frac{ 1 }{ 3*3*3 }=\frac{ 1 }{ 3*3*3 *3*3}=\frac{ 1 }{ 3^5 }\]

OpenStudy (kaleineedshelp):

@ememnim , yeah. It is 1 over 3 to the second power times 1 over 3 to the third power.

OpenStudy (kaleineedshelp):

@sshayer , is 3^-5 the same thing as 1/3^5 ?

OpenStudy (sshayer):

\[\frac{ 1 }{ 3^2 }=\frac{ 1 }{ 3*3 }\] yes\[\frac{ 1 }{ 3^5 }=3^{-5}\]

OpenStudy (kaleineedshelp):

@sshayer , I might have a couple other questions , can you help me with that as well or should I post another question for each one ?

OpenStudy (sshayer):

i will try.

OpenStudy (kaleineedshelp):

@sshayer (1/7^3)^4

OpenStudy (sshayer):

\[\left( \frac{ 1 }{ 7^3 } \right)^4=\frac{ 1^4 }{ 7^{3*4} }=\frac{ 1 }{ 7^{12} }\]

OpenStudy (sshayer):

\[\left( a^p \right)^q=a ^{pq}\]